OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles?

OpenStudy (anonymous):

what number is in the m position in y = -3x - 1/8

OpenStudy (anonymous):

2 in=25 mi.

OpenStudy (anonymous):

4.8 inches?

OpenStudy (anonymous):

is it 4.8 inches?

OpenStudy (anonymous):

no that will be 120 in

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

it is 2.4

OpenStudy (texaschic101):

2/25 = x/60 --- (2 in to 25 miles - x in to 60 miles) cross multiply 25(x) = 2(60) 25x = 120 -- divide both sides by 25 x = 120/25 x = 4.8 4.8 is the correct answer

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

How many miles are represented by 1 4 inch on the map? Enter your answer in the box. miles Save

OpenStudy (texaschic101):

is that 1/4 of an inch ??

OpenStudy (anonymous):

ya 1/4 inch

OpenStudy (anonymous):

so wats the ASNWER?

OpenStudy (texaschic101):

2/25 = (1/4)/x cross multiply 2(x) = 25(1/4) 2x = 25/4 -- divide both sides by 2 x = (25/4) / 2 x = 25/4 * 1/2 x = 25/8 = 3 1/8 miles

OpenStudy (anonymous):

I need help on that one

OpenStudy (texaschic101):

3 1/8 miles are represented by 1/4 inches

OpenStudy (anonymous):

whats the answer ?

OpenStudy (texaschic101):

or turn it onto a decimal.....1/4 inches = 3.125 miles or 3 1/8 miles

OpenStudy (texaschic101):

3 and 1/8 miles or 3.125 miles

OpenStudy (anonymous):

another question?

OpenStudy (anonymous):

Question 2: A circular picture is 8 inches in diameter. Part A What is the area of the picture in square inches? A. 4π square inches B. 8π square inches C. 16π square inches D. 32π square inches

OpenStudy (texaschic101):

A = 1/4 * pi * d^2 A = 1/4 * pi * 8^2 A = 1/4 * pi * 64 A = pi * 16.....so its 16 pi sq inches....its C